# Hybrid Control Technique for Switched LPV Systems and Its Application to Active Magnetic Bearing System

### v1

@@ 1 Introduction
Hybrid systems have received sustained attention in the control community due to their ability to model and control complex dynamical behaviors that cannot be adequately addressed by purely continuous or discrete frameworks. By integrating continuous dynamics with logical decision mechanisms, hybrid control architectures can effectively handle large uncertainties, nonlinearities, and mode-dependent behaviors, often alleviating fundamental limitations of conventional control designs and enabling improved performance and flexibility [7, 1, 15, 20]. As an important subclass of hybrid systems, switched linear systems which consisting of multiple linear subsystems and a rule governing switching among them, have been extensively studied [10, 12, 9]. Despite the availability of powerful linear analysis tools, stability and stabilization of switched systems remain challenging, since instability may arise even when all individual subsystems are asymptotically stable [11]. Consequently, the switching signal plays a critical role in determining system behavior and must be explicitly considered in controller design.
For autonomous switching, stabilization can be treated as a robust control problem, where the existence of a common Lyapunov function (CLF) guarantees stability under arbitrary switching [10]. However, CLF-based conditions are often overly conservative, particularly when specific switching logics are prescribed. This has motivated the development of multiple Lyapunov function (MLF) approaches [6, 28], which enable less conservative stability analysis under structured switching schemes such as dwell-time, average dwell-time, and hysteresis switching [17, 11]. Among these, hysteresis switching is particularly attractive as it implicitly enforces dwell-time, prevents chattering near switching boundaries, and improves robustness to noise and uncertainty.
While MLF-based methods have been successfully applied to state-feedback synthesis using LMI formulations, output-feedback control of switched systems remains substantially more difficult. In particular, enforcing boundary conditions between Lyapunov functions typically leads to nonconvex bilinear matrix inequalities (BMIs), which are computationally challenging. Existing approaches either decouple controller synthesis from switching conditions, resulting in conservative performance or excessive dwell-time requirements or rely on restrictive assumptions such as full state availability at switching instants [13]. More recently, hybrid output-feedback frameworks incorporating controller-state reset mechanisms have been developed for switched systems under dwell-time switching logic. This enabled convex LMI-based synthesis while explicitly accounting for switching-induced boundary conditions, significantly reducing conservatism and improving performance [29, 30].
@@ 2 Hybrid LPV Control Approach
After solving for the matrix functions $R_{i}(\rho)$ and $S_{i}(\rho)$, the controller gains are reconstructed using the standard formulas in [27]:
@@ 3.2 Hybrid LPV Controller Design and Simulation
Figure 5: LPV controller activition sequence.
@@ 4 Concluding Remarks
Future research will focus on extending the proposed framework to flexible-mode AMB models, where structural resonances, unmodeled dynamics, and uncertainty become increasingly important at high rotational speeds. The integration of robust and hybrid gain-scheduling techniques is expected to further enhance performance, stability margins, and reliability in advanced magnetic bearing applications.

### v2

@@ 1 Introduction
Hybrid systems have received sustained attention in the control community due to their ability to model and control complex dynamical behaviors that cannot be adequately addressed by purely continuous or discrete frameworks. By integrating continuous dynamics with logical decision mechanisms, hybrid control architectures can effectively handle large uncertainties, nonlinearities, and mode-dependent behaviors, often alleviating fundamental limitations of conventional control designs and enabling improved performance and flexibility [6, 1, 16, 21, 8]. As an important subclass of hybrid systems, switched linear systems which consisting of multiple linear subsystems and a rule governing switching among them, have been extensively studied [10, 25, 12]. Despite the availability of powerful linear analysis tools, stability and stabilization of switched systems remain challenging, since instability may arise even when all individual subsystems are asymptotically stable [11]. Consequently, the switching signal plays a critical role in determining system behavior and must be explicitly considered in controller design.
For autonomous switching, stabilization can be treated as a robust control problem, where the existence of a common Lyapunov function (CLF) guarantees stability under arbitrary switching [10]. However, CLF-based conditions are often overly conservative, particularly when specific switching logics are prescribed. This has motivated the development of multiple Lyapunov function (MLF) approaches [5, 29], which enable less conservative stability analysis under structured switching schemes such as dwell-time, average dwell-time, and hysteresis switching [18, 5, 29, 11]. Among these, hysteresis switching is particularly attractive as it implicitly enforces dwell-time, prevents chattering near switching boundaries, and improves robustness to noise and uncertainty.
While MLF-based methods have been successfully applied to state-feedback synthesis using LMI formulations, output-feedback control of switched systems remains substantially more difficult. In particular, enforcing boundary conditions between Lyapunov functions typically leads to nonconvex bilinear matrix inequalities (BMIs), which are computationally challenging. Existing approaches either decouple controller synthesis from switching conditions, resulting in conservative performance or excessive dwell-time requirements or rely on restrictive assumptions such as full state availability at switching instants [14, 13]. More recently, hybrid output-feedback frameworks incorporating controller-state reset mechanisms have been developed for switched systems under dwell-time switching logic. This enabled convex LMI-based synthesis while explicitly accounting for switching-induced boundary conditions, significantly reducing conservatism and improving performance [30, 31].
@@ 2 Hybrid LPV Control Approach
After solving for the matrix functions $R_{i}(\rho)$ and $S_{i}(\rho)$, the controller gains are reconstructed using the standard formulas in [28, 9]:
@@ 3.2 Hybrid LPV Controller Design and Simulation
Figure 5: LPV controller activation sequence.
@@ 4 Concluding Remarks
Future research will focus on extending the proposed framework to flexible-mode AMB models, where structural resonances, unmodeled dynamics, and uncertainty become increasingly important at high rotational speeds [3]. The integration of robust and hybrid gain-scheduling techniques is expected to further enhance performance, stability margins, and reliability in advanced magnetic bearing applications.